what is the unknown of log5 v + log1/3 9 + log 9 81 = -4
Answer by stanbon(75887)

what is the unknown of log5 (v) + log1/3 (9) + log 9 (81) = -4
-------------------------
Note:
Solve:
(1/3)^x=9
3^-x = 3^2
x = -2
------
Solve:
9^x = 81
8^x = 9^2
x = 2
----
Put those values where they belong and solve for "v":
log5(v) + (-2) + (2) = -4
---
log5(v) = -4
v = 5^(-4)
---
v = 1/625
===================
Cheers,
Stan H.
===================